OpenStudy (anonymous):

Pam likes to practice dancing while preparing for a math tournament. She spends 80 minutes every day practicing dance and math. To help her concentrate better on math, she spends 20 minutes more dancing than doing math.

OpenStudy (anonymous):

Part A: Write a pair of linear equations to show the relationship between the number of minutes Pam practices math every day (x) and the number of minutes she practices dances every day (y). Part B: How much time does Pam spend on practicing math every day? Part C: Is it possible for Pam to have spent 60 minutes practicing dance every day? Explain your reasoning.

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

What is the eqn for this "80 minutes every day practicing dance and math"?

OpenStudy (superdavesuper):

And then what is the eqn for "20 minutes more dancing than doing math"?

OpenStudy (anonymous):

idk ...

OpenStudy (superdavesuper):

Just write it out in terms of x and y - it is not difficult :) trust me!

OpenStudy (anonymous):

@superdavesuper x=20 ?

OpenStudy (superdavesuper):

I will do the first one for u: x+y=80....how about the second eqn?

OpenStudy (anonymous):

x+20=80 ???

OpenStudy (superdavesuper):

Nope....it is something like y-x=...

OpenStudy (anonymous):

y-x=20 ?

OpenStudy (superdavesuper):

:) yes so these are the answers for Part A

OpenStudy (anonymous):

ok cool... so part b

OpenStudy (superdavesuper):

Part b is solving for x...subtract the 1st eqn from the 2nd...what do u get?

OpenStudy (anonymous):

wait so y-x=20 - x+y=80 ???

OpenStudy (superdavesuper):

try again but left side stays on the left side of the =....same for the right side....

OpenStudy (anonymous):

idk !! >.<

OpenStudy (anonymous):

ok so x=30 and y =50 .. now what ?

OpenStudy (superdavesuper):

On to Part C of course :)

OpenStudy (anonymous):

wait whats part b

OpenStudy (superdavesuper):

But u got the answer to dat already....cuz' u got the solution for y..

OpenStudy (superdavesuper):

Part b asks u "How much time does Pam spend on practicing math every day?"

OpenStudy (anonymous):

50 minutes ? she spends an extra 20 on dancing and that only equals 70

OpenStudy (superdavesuper):

No, number of minutes Pam practices math every day (x)

OpenStudy (anonymous):

its 50 mins?

OpenStudy (superdavesuper):

Nope what is X? u solved it earlier...scroll up n see plz

OpenStudy (anonymous):

ohhh wtf, 30

OpenStudy (anonymous):

x=30 and y=50

OpenStudy (superdavesuper):

Thats Part B :)

OpenStudy (anonymous):

so she spends 30 mins on math?

OpenStudy (superdavesuper):

Yup so how much time on dance? ;)

OpenStudy (anonymous):

and this is my response for c Yes because she spends a total of 80 minutes. However, she would have to reduce her math time by 10 minutes, leaving only 20 minutes for math and 60 minutes for dance.

OpenStudy (anonymous):

dance is 50 mins

OpenStudy (superdavesuper):

You got it :)

OpenStudy (anonymous):

yaayyyy

OpenStudy (anonymous):

ok ready for another question ?

OpenStudy (superdavesuper):

OK u have to hurry as i have 10 min left..

OpenStudy (anonymous):

A system of equations is shown below. 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)